a rectangular field has a total perimeter of 144 feet the width is 36 feet less than the lenght.
Answer by TimothyLamb(4379)

2L + 2w = 144
w = 36
---
2L + 2w = 144
2L + 2(36) = 144
2L + 72 = 144
2L = 72
---
L = 36
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php